OpenStudy (anonymous):

evaluate with l'hopital's rule the limit as x goes to 0 of xcscx

hartnn (hartnn):

@ericalaine Hi, \(\Huge \mathcal{\text{Welcome To OpenStudy}\ddot\smile} \) you can write csc x = 1/ sin x to be able to apply L'Hopital's rule.

OpenStudy (anonymous):

would it not be cscx/(1/x) then apply rules because that is what my book told me to do but then I'm stuck after that.....

hartnn (hartnn):

you can do that also but that would complicate things. because if you differentiate 1/x, you'll get -1/x^2. and then you have to take that -x^2 in the denominator to get -x^2 *(-cot^2 x) and then if you put x=0, you get the answer.

hartnn (hartnn):

whereas, if you do, x/ sin x after applying L'Hopital's you directly get 1/ cos x quite simple, right ?

OpenStudy (anonymous):

yes way more simple!

hartnn (hartnn):

:) any more doubts ?

OpenStudy (anonymous):

where does sin come from exactly?

hartnn (hartnn):

oh, didn't you know, \(\large \sin x=\dfrac{1}{\csc x}\)

OpenStudy (anonymous):

thank you so much!

hartnn (hartnn):

welcome ^_^

hartnn (hartnn):

oh, and since you are new here, if you have any doubts browsing this site, you can ask me, or to an ambassador (with an encircled A before their name). Have fun learning with Open Study! :)